OpenStudy (anonymous):

Will fan and medal! Solve. 2x2 + x – 1 = 2

OpenStudy (anonymous):

@Hero

OpenStudy (anonymous):

try it by using fectorization method

OpenStudy (anonymous):

the above equation will become 2x2 + x -1 =2 2x2 + x -1 - 2 = 0 2x2 + x - 3 = 0

OpenStudy (triciaal):

then from the form ax^2 + bx + c = 0 find your AC what are the factors of AC that add up to B , pay attention to the signs

OpenStudy (triciaal):

using -2 and +3 2x^2 -2x + 3x -3 = 0 then group and factor 2x(x -1) +3(x-1)= 0 use the distributive property with the common factor (2x + 3)( x - 1) = 0 to get a product of zero then one factor must be zero 2x - 3 = 0 or x - 1 = 0